# Computable conditions for order-2 $CP$ symmetry in NHDM potentials

### v1

@@ 3.3 Numerical considerations
Some comments about the practical implementation of $CP2$ detection by means of Algorithms 1 and 2 are in order. First, it is often the case (e.g. in a uniform parameter scan) that a symmetry cannot truly be exact. This can be due to the symmetric subset of the parameter space having measure zero, or simply finite numerical precision. Either way, an appropriate numerical tolerance has to be defined such that parameter points which are sufficiently close to exact symmetry are considered symmetric. We want to emphasize that Algorithms 1 and 2 can be implemented with such a tolerance in order to detect parameter points close to exact $CP2$ symmetry. Indeed, if the tolerance is encoded by a small number $\epsilon$, then it suffices to neglect all numbers smaller than $\epsilon$ in the numerical computations.
Secondly, even though large values of $N$ have limited practical applications to e.g. phenomenology, one might wonder about the computational cost of Algorithm 2 and how it scales with $N$. The most expensive step is checking for a subalgebra since that requires the computation of the structure constants (3.1) [^4] With the exception of the pathological cases discussed in 3.2.1 for which the required number of operations scales as $N^{12}$. While the computation time increases fast, the presence or absence of $CP2$ can be established almost instantaneously for $N=3$ and $N=4$ doublets which are, arguably, the most important use cases.
@@ 4.1 $N=3:$ The Ivanov-Silva potential
showing that the representation has two highest weights with Dynkin labels 0 and 1. The potential given by (4.9)-(4.14) has therefore, as expected, no $CP2$ symmetry since $\Lambda$ has three eigenvectors forming a basis for the representation $\mathbf{1+2}$ of ${\mathfrak{so}}(3)$ while it is $\mathbf{3}$ which corresponds to $CP2$. Actually the representation $\mathbf{1+2}$, accompanied by the $LM$-orthogonality conditions, corresponds to a different block structure which partially characterizes $CP4$ in 3HDMs [12].
@@ 4.2 $N=4:\>\mathbb{Z}_{6}$-symmetric potential
$m^{2}$ $=1,\quad\Lambda=1,\quad\Lambda^{\prime}=2,\quad\Lambda^{\prime\prime}=3,\quad\tilde{\Lambda}^{\prime}=4,\quad\tilde{\Lambda}^{\prime\prime}=-\frac{1}{2},\quad\lambda_{1}=i,$ (4.39)
$\lambda_{2}$ $=i,\quad\lambda_{3}=e^{i\frac{2\pi}{3}},\quad\lambda_{4}=1,\quad\lambda_{5}=2.\quad$ (4.40)
@@
Since $L=M=0$, the orthogonality conditions are automatically satisfied by all eigenvectors of  (4.36) and we proceed to compute the ${\mathfrak{su}}(N)$ structure constants $Z_{(ij)k}$ in the basis of eigenvectors, in order to look for 6d subalgebras. Due to its large size, displaying the $Z$-matrix is impractical and not very illuminating, therefore we omit it. Nevertheless, the subalgebra search is easily done using a computer program to implement the strategies explained in Sect. 4.1, and it is found that two sets of eigenvectors generate 6d subalgebras of rank 2, for which the only candidate is ${\mathfrak{so}}(4)$. Thus it is not necessary to compute the root system, and it only remains to identify which 4d representations have been found. From Proposition 2, the only possibilities are $\mathbf{4}$, i.e. the defining representation, and the reducible representation $\mathbf{2+2^{\prime}}$. Computing the Dynkin labels of both representations as described in Sect. 2.2 one finds
@@ 4.3 $N=7$
As a last example which, while mostly academic, shows the power of this method for $CP2$ detection, we apply Algorithm 3 to a 7HDM potential whose parameter values are given in Appendix B. From a Lie algebraic perspective, $N=7$ is interesting as it is the first value where there exists a semi-simple Lie algebra with the same dimension and rank as ${\mathfrak{so}}(N)$, but which is not isomorphic to it, namely $\mathfrak{sp}(6)$.
(a) ${\mathfrak{so}}(7)$
(b) $\mathfrak{sp}(6)$
Algorithm 1 reveals that a maximal set of orthonormal eigenvectors satisfying $LM$-orthogonality has 41 elements. With so many candidate eigenvectors, searching for the subalgebra pattern (3.2) in ${\mathfrak{su}}(7)$ starts to become computationally expensive. For reference, running our implementation of Algorithm 3 on an ordinary computer it takes less than a minute to find that 21 eigenvectors close under the F-product, forming a 21d rank-3 subalgebra. The root systems of the two possible algebras, ${\mathfrak{so}}(7)$ and $\mathfrak{sp}(6)$, which differ only by the lengths of the roots, are shown in Fig. 1. In the example at hand one finds that the root system of the algebra to be identified is in fact that of $\mathfrak{sp}(6)$, and hence the corresponding potential has no $CP2$ symmetry.
@@ 5 Summary
We have derived necessary and sufficient conditions for an NHDM potential to admit a $CP2$ symmetry, which are formulated as relations among vectors that transform according to the adjoint representation under an $SU(N)$ change of doublet basis. Such vectors can naturally be thought of elements of ${\mathfrak{su}}(N)$, which allows one to use the Lie algebra structure to verify basis-invariant properties such as being related to a particular subspace in the adjoint space. In the case of $CP2$, the relevant subspace actually corresponds to a ${\mathfrak{so}}(N)$ subalgebra of ${\mathfrak{su}}(N)$ and the main task for detecting this symmetry is checking whether a subset of eigenvectors of the bilinear quadratic form $\Lambda$ generates the defining representation of ${\mathfrak{so}}(N)$. By considering all the $k$-dimensional subalgebras of ${\mathfrak{su}}(N)$ and all the $N$-dimensional representations of ${\mathfrak{so}}(N)$ we developed an optimized computable algorithm for this task. The complete algorithm for detecting $CP2$ works in principle for any number of doublets $N$, and is only limited by computational cost. We find that, running our algorithm on a regular desktop computer, a generic parameter space point can be labelled $CP2$-conserving or $CP2$-violating in less than a minute for $N\leq 7$. However, when a $CP2$ symmetry exists, finding a real basis explicitly in general remains out of reach.
@@ Appendix B Parameter values for the $N=7$ numerical example
Below are the numerical values for the parameter point used in the example analyzed in Section. 4.3. All the non-zero elements are listed, except those which can be obtained by symmetry of $\Lambda$.

### v2

@@ 3.3 Numerical considerations
Some comments about the practical implementation of $CP2$ detection by means of Algorithms 1 and 2 are in order. First, all the steps in these algorithms are linear algebra computations which can, in principle, be carried out analytically. However, a complete analytic treatment would require very simple expressions for the eigenvectors of $\Lambda$ which is unlikely to be the case for non-trivial potentials. Thus, in practice, a numerical implementation of the algorithms is most relevant.
Secondly, it is often the case (e.g. in a uniform parameter scan) that a symmetry cannot truly be exact. This can be due to the symmetric subset of the parameter space having measure zero, or simply finite numerical precision. Either way, an appropriate numerical tolerance has to be defined such that parameter points which are sufficiently close to exact symmetry are considered symmetric. We want to emphasize that Algorithms 1 and 2 can be implemented with such a tolerance in order to detect parameter points close to exact $CP2$ symmetry. Indeed, if the tolerance is encoded by a small number $\epsilon$, then it suffices to neglect all numbers smaller than $\epsilon$ in the numerical computations.
Lastly, even though large values of $N$ have limited practical applications to e.g. phenomenology, one might wonder about the computational cost of Algorithm 2 and how it scales with $N$. The most expensive step is checking for a subalgebra since that requires the computation of the structure constants (3.1) [^4] With the exception of the pathological cases discussed in 3.2.1. for which the required number of operations scales as $N^{12}$. While the computation time increases fast, the presence or absence of $CP2$ can be established almost instantaneously for $N=3$ and $N=4$ doublets which are, arguably, the most important use cases.
@@ 4.1 $N=3:$ The Ivanov-Silva potential
showing that the representation has two highest weights with Dynkin labels 0 and 1. The potential given by (4.9)–(4.14) has therefore, as expected, no $CP2$ symmetry since $\Lambda$ has three eigenvectors forming a basis for the representation $\mathbf{2+1}$ of ${\mathfrak{so}}(3)$ while it is $\mathbf{3}$ which corresponds to $CP2$. Actually the representation $\mathbf{2+1}$, accompanied by the $LM$-orthogonality conditions, corresponds to a different block structure which partially characterizes $CP4$ in 3HDMs [13].
@@ 4.2 $N=4:\>\mathbb{Z}_{6}$-symmetric potential
$m^{2}$ $=1,\quad\Lambda=1,\quad\Lambda^{\prime}=2,\quad\Lambda^{\prime\prime}=3,\quad\tilde{\Lambda}^{\prime}=4,\quad\tilde{\Lambda}^{\prime\prime}=-\frac{1}{2},\quad\lambda_{1}=i,$
$\lambda_{2}$ $=i,\quad\lambda_{3}=e^{i\frac{2\pi}{3}},\quad\lambda_{4}=1,\quad\lambda_{5}=2,\quad$ (4.39)
@@
Since $L=M=0$, the orthogonality conditions are automatically satisfied by all eigenvectors of (4.36) and we proceed to compute the ${\mathfrak{su}}(N)$ structure constants $Z_{(ij)k}$ in the basis of eigenvectors, in order to look for 6d subalgebras. Due to its large size, displaying the $Z$-matrix is impractical and not very illuminating, therefore we omit it. Nevertheless, the subalgebra search is easily done using a computer program to implement the strategies explained in section 4.1, and it is found that two sets of eigenvectors generate 6d subalgebras of rank 2, for which the only candidate is ${\mathfrak{so}}(4)$. Thus it is not necessary to compute the root system, and it only remains to identify which 4d representations have been found. From Proposition 2, the only possibilities are $\mathbf{4}$, i.e. the defining representation, and the reducible representation $\mathbf{2+2^{\prime}}$. Computing the Dynkin labels of both representations as described in section 2.2 one finds
@@ 4.3 $N=7$
As a last example which, while mostly academic, shows the power of this method for $CP2$ detection, we apply Algorithm 3 to a 7HDM potential whose parameter values are given in appendix B. From a Lie algebraic perspective, $N=7$ is interesting as it is the first value where there exists a semisimple Lie algebra with the same dimension and rank as ${\mathfrak{so}}(N)$, but which is not isomorphic to it, namely $\mathfrak{sp}(6)$.
Algorithm 1 reveals that a maximal set of orthonormal eigenvectors satisfying $LM$-orthogonality has 41 elements. With so many candidate eigenvectors, searching for the subalgebra pattern (3.2) in ${\mathfrak{su}}(7)$ starts to become computationally expensive. For reference, running our implementation of Algorithm 3 on an ordinary computer it takes less than a minute to find that 21 eigenvectors close under the F-product, forming a 21d rank-3 subalgebra. The root systems of the two possible algebras, ${\mathfrak{so}}(7)$ and $\mathfrak{sp}(6)$, which differ only by the lengths of the roots, are shown in figure 1. In the example at hand one finds that the root system of the algebra to be identified is in fact that of $\mathfrak{sp}(6)$, and hence the corresponding potential has no $CP2$ symmetry.
@@ 5 Summary
We have derived necessary and sufficient conditions for an NHDM potential to admit a $CP2$ symmetry, which are formulated as relations among vectors that transform according to the adjoint representation under an ${\mathsf{SU}}(N)$ change of doublet basis. Such vectors can naturally be thought of elements of ${\mathfrak{su}}(N)$, which allows one to use the Lie algebra structure to verify basis-invariant properties such as being related to a particular subspace in the adjoint space. In the case of $CP2$, the relevant subspace actually corresponds to a ${\mathfrak{so}}(N)$ subalgebra of ${\mathfrak{su}}(N)$ and the main task for detecting this symmetry is checking whether a subset of eigenvectors of the bilinear quadratic form $\Lambda$ generates the defining representation of ${\mathfrak{so}}(N)$. By considering all the $k$-dimensional subalgebras of ${\mathfrak{su}}(N)$ and all the $N$-dimensional representations of ${\mathfrak{so}}(N)$ we developed an optimized computable algorithm for this task. The complete algorithm for detecting $CP2$ works in principle for any number of doublets $N$, and is only limited by computational cost. We find that, running our algorithm on a regular desktop computer, a generic parameter space point can be labelled $CP2$-conserving or $CP2$-violating in less than a minute for $N\leq 7$. However, when a $CP2$ symmetry exists, finding a real basis explicitly in general remains out of reach.
@@ Appendix B Parameter values for the $N=7$ numerical example
Below are the numerical values for the parameter point used in the example analyzed in section 4.3. All the non-zero elements are listed, except those which can be obtained by symmetry of $\Lambda$.